OpenStudy (anonymous):

which ratio forms a proportion with 6/14 A.12/7 B.12/28 C.7/3 D.4/12

OpenStudy (blurbendy):

6/14 = 3/7 12/28 = 3/7

OpenStudy (anonymous):

@blurbendy thank you :)

OpenStudy (blurbendy):

np